OpenStudy (anonymous):

Complete the equation for the linear function whose graph contains the point (7, 8) and has a slope of −1.2. y =____x +____ @freckles

OpenStudy (freckles):

@TeenWolfGirl can i see your attempt on this one

OpenStudy (freckles):

You have to test what we have gone over at some point.

OpenStudy (freckles):

I think while I'm here is the best time to do that.

OpenStudy (anonymous):

ok i am back well it tells me my slope it -1.2

OpenStudy (stamp):

Your slope m is -1.2, so knowing y = mx + b and knowing the slope m, you can begin filling out your equation. y = -1.2x + b

OpenStudy (anonymous):

y - 8 = -1.2* ( x - 7) y - 8 = -1.2x + 8.4 y = -1.2x + 16.4

OpenStudy (anonymous):

correct?

OpenStudy (stamp):

In math, you use the things you know and put them together to help find what you do not know. Knowing the slope m and knowing y = mx + b allows you to put those two thoughts together to move closer toward your objective.

OpenStudy (stamp):

y - 8 = -6/5 (x - 7) y = -6/5 x + 42/5 + 40/5 y = -6/5x + 82/5 y = -1.2x + 16.4

OpenStudy (stamp):

You are correct, well done.

OpenStudy (freckles):

good job

OpenStudy (anonymous):

thanks